OpenStudy (anonymous):

Find value without calculator or table (in radians) csc^-1 (2/squareroot2), sec^-1 (2/squareroot2)

OpenStudy (anonymous):

FOr this type of question, you have to remember the unit circle.When converting these back to sin and cos, the ordered pair if \[(\frac{\sqrt{2}}{2}, \frac{\sqrt{2}}{2})\]What angle measure on the unit circle has this ordered pair?

OpenStudy (anonymous):

pi/4

OpenStudy (anonymous):

never mind..dont know

OpenStudy (anonymous):

\[\huge\text{You got it! }\frac{π}{4}\text{ is correct!}\]

OpenStudy (anonymous):

yeah but how do I show work on test??

OpenStudy (anonymous):

You can't really show much work...if you want to show work, you couls show how sec and csc are reciprocals of sin and cos which is how I got the √2/2 stuff. Then you could draw the first quadrant of the unit circle including the value of π/4.

OpenStudy (anonymous):

Is there a trick for memorizing unit circle??

OpenStudy (anonymous):

Well really, you just have to focus on the first quadrant. The using logic, you can remember the other three using the first. Yeah. The two values I showed are the same because they are reflected, but since it's in quadrant 2, the x-value is negative.

OpenStudy (anonymous):

Also remember All Students Take Calculus

OpenStudy (anonymous):

It represents the values out of sin, cos, and tan are positive in each quadrant. All (Quadrant I) Sin (Quadrant II) Tan (Quadrant III) Cos (Quadrant IV) See?

OpenStudy (anonymous):

Yep thanks

OpenStudy (anonymous):

np :)